a) A university wishes to estimate the average number of hours students spend doing homework per week. The variance obtained from previous study is 25 hours. How large a sample must be selected if they want to be 98% confident of finding the true mean given the error is within 1.5 hours?
(b) Find t value for a 95% confidence interval when the sample size is 8.
1
Expert's answer
2020-10-26T19:09:18-0400

a) If "\sigma" not known:

If "n" is large, then use normal.

If "n" is small, then use t-distribution.

Margin of error


"ME=t_{\alpha/2, df}\dfrac{s}{\sqrt{n}}"

Given "ME=1.5\ h, s=25\ h, \alpha=0.02"

If "n\leq30"


"ME=t_{\alpha/2, df}\dfrac{25}{\sqrt{n}}>1/5"

Hence we can use normal distribution


"ME=z_{\alpha/2}\dfrac{s}{\sqrt{n}}"


"z_{\alpha/2}\dfrac{25}{\sqrt{n}}\leq1.5"

"n\geq(\dfrac{25z_{\alpha/2}}{1.5})^2"

"z_{0.02\/2}=2.3263"

"n\geq(\dfrac{25\cdot2.3263}{1.5})^2"

"n\geq1504"



(b)

"t_{\alpha/2, df}=t_{0.05/2, 8-1}=2.364619"